# Direct probe of magnetic field effects on phonons by ultrasound propagation in a quasi-two-dimensional honeycomb magnet Na2Co2TeO6

### v1

@@ Abstract
We study the phonon behavior of a Co-based honeycomb frustrated magnet Na2Co2TeO6 under magnetic field applied perpendicular to the honeycomb plane. The temperature and field dependence of the sound velocity and sound attenuation unveil prominent spin-lattice coupling in this material, promoting ultrasound as a sensitive probe for magnetic properties. An out-of-plane ferrimagnetic order is determined below the Néel temperature $T_{N}=27$ K. A comprehensive analysis of our data further provides strong evidence for a triple-Q ground state of Na2Co2TeO6. Furthermore, the ultrasound data underscore that the field impact on the thermal conductivity as recently reported is of pure phononic origin.
@@ I Introduction
It is fair to say that renewed attention to the quasi-two-dimensional cobalt systems was drawn by the proposal that Co-based honeycomb magnets might realize Kitaev QSL physics Liu2018 ; Sano2018 ; Liu2020 . A wave of experiments have revealed many resemblances between Na2Co2TeO6 (NCTO) and the leading Kiteav QSL candidate $\alpha-$RuCl3 Yao2020 ; Lin2021 ; Hong2021 ; Sanders2022 ; Yang2022 ; Gillig2023 ; Chen2024 ; Zhou2024 . However, NCTO exhibits an out-of-plane ferrimagnetic order Yao2020 , which is incompatible with a pure Kitaev QSL ground state. Furthermore, detailed theoretical calculations suggested that the Kitaev interaction, while present, is not the dominant term in NCTO’s Hamiltonian Winter2022 . These facts fuel the debates about whether the exotic thermal transport results reported in $\alpha-$RuCl3, once regarded as signatures of QSL states, are just from phonons influenced by magnetic excitation Hentrich2018 ; Lefrancois2022 ; Bruin2022 . Practically, extracting the putative charge-neutral electronic excitations of interest from a phonon background invokes comparison with a nonmagnetic reference compound or a power-law fitting of the $\kappa_{ph}(T)$ curves in a limited temperature range. Both of these procedures are tricky and proliferate the discrepancies in interpreting the data. In this context, it is important to get more detailed knowledge of the phonons in this material class.
@@
In this work, we focus on measuring the impact of out-of-plane ($\bf B$//$\bf c$) magnetic fields on the in-plane ultrasound properties of NCTO. Our results demonstrate that ultrasound serves as a sensitive probe of structural and magnetic transitions in NCTO. Very rich phenomena are revealed below its Néel temperature ($T_{N}=27$ K). By analyzing our results, we confirm the existence of an out-of-plane ferrimagnetic order in NCTO and provide evidence supporting a triple-Q ground state Kruger2023 ; Chen2021 . Furthermore, we explore the relationship between our ultrasound data and the reported thermal transport results under out-of-plane magnetic fields Yang2022 ; Hong2024 , offering new insights into the interplay between lattice dynamics and magnetic excitations in this material and beyond.
Figure 2: (a) Field dependence of $\Delta v/v_{0}$ at different temperatures. The curves are shifted vertically for clarity. Note the different rescale factors for each curve. The black arrows indicate the dips in some curves. (b) The gray circles, squares, and dashed lines are phase boundaries of NCTO inferred from its magnetic susceptibility data Zhang2024b . The red stars are the positions of anomalies from Fig. 2(a). We follow the notation for the critical fields $H_{1}$ and $H_{2}$ from Ref. Zhang2024b . Although there is a clear line below $H_{1}$ in the contour plot of $dM/dH$ data, it was not labeled as a critical field in Ref. Zhang2024b . Nevertheless, an electron spin resonance (ESR) work identified it also as a phase transition Bera2023 . Thus we use $H_{0}$ to name this line. The dashed black lines indicate the parameter space reached in Fig. 2(a).
@@ II Experimental Method
Single crystals of NCTO were prepared with a modified flux method Yao2020 . Although the typical thickness of the as-grown crystals are about 0.1 mm, very few of them can be thicker, preferred by an ultrasound experiment Yao2020 . Three of the thickest crystals, all with a thickness of 0.4$\pm$0.05 mm, were selected from different batches for this work. NCTO has a hexagonal crystal structure (space group $P$6322) Zhang2023 . The ultrasound propagation direction with the wave vector $k$ was set along the crystallographic zigzag $\bf a$-direction (perpendicular to a Co-Co bond in the honeycomb plane, see the insets of Fig. 1) for simplicity. The polarization of the acoustic wave ($u$) was along $\bf a$, $\bf b$ or $\bf c$ crystallographic directions depending on the studied acoustic mode. The direction $\bf b$ is along Co-Co bond within the honeycomb plane. The surfaces to which the transducers would be attached were carefully polished against ultra-fine grade abrasive papers.
@@ III Results and Discussion
The temperature dependence of the relative changes in the sound velocity $\Delta v/v_{0}$ for the longitudinal acoustic (LA) mode is shown in Fig. 1(a). The overall $\Delta v/v_{0}(T)$ curve reveals the lattice hardening upon cooling, and a softening appears below the magnetic transition temperature $T_{N}$. The data at low temperature is enlarged in Fig. 1(b), plotted together with results taken with various ultrasound frequencies, to check for a possible frequency dependence. All these curves overlap almost perfectly. Three anomalies can be reliability resolved, corresponding to the cascade of magnetic phase transitions below $T_{N}$ reported in the literature Lefrancois2016 ; Bera2017 .
Figure 3: Field dependence of (a) $\Delta v/v_{0}$ and (b) $\Delta\alpha$ at 1.8 K, measured in opposite field directions. The first field-sweep run was done after the sample was cooled in zero, or in small ($<14$ mT) remanent field of our magnet. The second and third measurements were done after cooling in a finite field applied along opposite directions.
@@
A set of field-dependent $\Delta v/v_{0}$ isotherms are presented in Fig. 2(a). Above 5 K, the $\Delta v/v_{0}(H)$ curves show moderate lattice softening in field without additional features up to $7.5$ T. The amplitude of changes in $\Delta v/v_{0}(H)$ increases considerably upon cooling. A clear dip in $\Delta v/v_{0}(H)$ can be resolved at 3 K. This dip gradually shifts to a lower field at lower temperatures. At the lowest temperature of 32 mK another broader dip at round 16 T was detected. These features are summarized in Fig. 2(b) in the phase diagram of NCTO for an out-of-plane magnetic field, together with the transitions observed in the magnetic susceptibility Zhang2024b . While all the ultrasound anomalies can find their correspondence in magnetization, some magnetic transitions leave no detectable fingerprint in our ultrasound data. This fact means that the acoustic modes in NCTO are selectively coupled to the spins. For the studied acoustic modes, the magnetoelastic coupling can be very small at some phase transitions in NCTO.
Figure 4: (a) $\Delta v/v_{0}(H)$ isotherms measured in NCTO in opposite field directions. The curves are shifted vertically for clarity. (b) The temperature dependence of the integrated hysteretic region of $\Delta v/v_{0}(H)$ isotherms. The data integrated up to 7.5 T from different experimental setting are rescaled to show a common temperature behavior. The yellow stripe is a guild to the eye. (c) Schematic diagram of the temperature and magnetic field effects on the out-of-plane magnetic components of two sub-lattices (red and blue arrows) with opposite sign and different magnitude, resides on a representative hexagon. Around the temperature at which the the energy of the opposite configurations in magnetic field is reversed, magnetic flipping is more likely to be promoted.
Figure 5: Field dependence of $\Delta v/v_{0}$ and $\Delta\alpha$ at (a) 630 mK, and (c) 15 K. The corresponding estimates of the change of phonon conductivity (see text) are plotted in red in (b) and (d), respectively. Field dependence of the thermal conductivity $\kappa$ at similar temperatures reported in the literature is also shown Yang2022 ; Hong2024 . Temperature dependence of maximum field-induced changes of (e) $\Delta v/v_{0}$ and (f) $\Delta\alpha$ measured with two different frequencies.
@@
Recently, there have been some debates on whether the magnetic ground state of NCTO is an unusual triple-Q state or the conventional multi-domain zigzag order Chen2021 ; Yao2023 ; Jin2025 ; Zhang2023 ; Dhanasekhar2025 . To shed more light on this issue, we investigated field dependence of $\Delta v/v_{0}$ and $\Delta\alpha$ under different cooling conditions (Fig. 3). We observe a very clear field-asymmetry depending on the cooling conditions, and it can be switched by changing the field cooling process from above $T_{N}$. This is a clear evidence of an out-of-plane ferromagnetic moment. It is compatible with the reported out-of-plane ferrimagnetic behavior in NCTO Yao2020 . There are two mechanisms proposed for the ferrimagnism in NCTO, both invoking the canting of the in-plane magnetic moments. Some previous work suggested a triple-Q ground state in NCTO and several other Co-based quasi-2D honeycomb magnets Chen2021 ; Yao2023 ; Jin2025 ; Kruger2023 ; Francini2024 . Another scenario insists that many cobaltates manifest a zigzag ground state stabilized within an XXZ model Zhang2023 ; Zhang2024b ; Dhanasekhar2025 ; Winter2022 , and the out-of-plane ferrimagnetism is due to unbalanced domains. Note that the zero-field-cooled (ZFC) $\Delta v/v_{0}(H)$ and $\Delta\alpha(H)$ data perfectly match the field-cooled (FC) data of the fully-switched states. It is unlikely that the alignment of domains can be made by a small remanent field. Hence, a more natural explanation is provided by a domain-free triple-Q ground state of NCTO.
@@
To gain more insight, additional $\Delta v/v_{0}(H)$ results at elevated temperatures all the way up to above $T_{N}$ are shown in Fig. 4(a). Obviously, the $\Delta v/v_{0}(H)$ curves are weakly field-suppressed and quite symmetric at temperatures higher than $T_{N}$. A butterfly-shaped hysteresis starts to grow at low temperatures. The hysteretic region and asymmetry are getting larger upon lowering the temperature. Finally, the hysteresis quickly shrinks, leaving very asymmetric $\Delta v/v_{0}(H)$ curves below about 12 K.
We use the integral of the enclosed region of the $\Delta v/v_{0}(H)$ curves to quantify the degree of hysteresis [Fig. 4(b)]. Data collected with different experimental settings were rescaled to fit into the same figure and to show a common temperature behavior for different acoustic modes. Notably, the hysteresis grows rapidly below $T_{N}$, reaches a maximum at around 12 K, and then drops quickly toward zero at lower temperature. As illustrated in Fig. 4(c), this behavior indicates that the energy difference between two magnetic configurations shrinks at around 12 K, not compatible with a XXZ ground state with domains developing below $T_{N}$. On the other hand, a recent Monte Carlo simulation work deduced a compensation point $T_{cp}$ at roughly $T_{N}/2$ for a sign reversion of the residual magnetization out of the honeycomb plane, if a triple-Q model is adopted for NCTO Francini2024 .
@@
$\kappa\propto Cvl=Cv^{2}\tau\propto\frac{Cv^{2}}{\Gamma},$ (1)
@@
$\Delta\kappa\propto\frac{C(v_{0}+\Delta v)^{2}}{\Gamma_{0}+\Delta\Gamma}-\frac{Cv_{0}^{2}}{\Gamma_{0}}.$ (2)
@@
$\frac{\Delta\kappa}{\kappa}\approx 2\times\frac{\Delta v}{v_{0}}-\frac{\Delta\Gamma}{\Gamma_{0}+\Delta\Gamma},$ (3)
@@
$\Delta\alpha\equiv-\frac{20}{L}\log_{10}\frac{I}{I_{0}},$ (4)
@@
$\frac{\Gamma_{0}}{\Gamma}=\frac{\tau}{\tau_{0}}\propto\frac{I}{I_{0}}=10^{-\Delta\alpha\times L/20}.$ (5)
@@
$\frac{\Delta\kappa}{\kappa}\approx 2\times\frac{\Delta v}{v_{0}}+10^{-\Delta\alpha\times L/20}-1.$ (6)
The estimated relative change in phonon thermal conductivity based on ultrasound data is shown in Fig. 5, by taking the sample thickness $L=0.4$ mm in Eq. (6). The effect of the out-of-plane magnetic field on the thermal transport of NCTO at nearly the same temperature reported in the literature is also sketched for comparison Yang2022 ; Hong2024 . Their shape roughly matches, but the amplitude calculated from phonon data is underestimated by an order of magnitude or more. The reason might be that only one acoustic mode at a single frequency is considered in this work, while all phonon modes contribute to phonon thermal transport. As indicated in Fig. 5(f), the $\Delta\alpha$ term is strongly frequency dependent. Nevertheless, both $\Delta v/v$ and $\Delta\alpha$ peak at near 3 K, a temperature range that has not been investigated in previous work. We expect an impressively large field suppression of $\kappa$ in this temperature range.
We are aware of the lowest temperature $\Delta v/v_{0}(H)$ curve reported in a recent ultrasound work on $\alpha-$RuCl3 features multiple humps and dips above its critical fields Hauspurg2024 . In the context of our results and discussion, our finding adds weight to the argument that the oscillatory structures in $\kappa(H)$ of $\alpha-$RuCl3 Czajka2021 , initially attributed to some charge neutral magnetic excitations, might actually come from the strongly scattered phonon transport channels Bruin2022 ; Lefrancois2023 .

### v2

@@ Abstract
We study the phonon behavior of a Co-based honeycomb frustrated magnet Na2Co2TeO6 under magnetic field applied perpendicular to the honeycomb plane. The temperature and field dependence of the sound velocity and sound attenuation unveil prominent spin-lattice coupling in this material, promoting ultrasound as a sensitive probe for magnetic properties. An out-of-plane ferrimagnetic order is determined below the Néel temperature $T_{N}=27$ K. A comprehensive analysis of our data further supports a triple-Q ground state of Na2Co2TeO6. Furthermore, the ultrasound data were systematically compared to the thermal transport results from literature, to unveil the importance of phononic contribution to the observed transport behaviors.
@@ I Introduction
However, the pervasiveness of phonons also poses significant challenges for reliably extracting tantalizing unconventional quasiparticle excitations. For instance, thermal conductivity ($\kappa$) at low temperatures, a powerful probe for charge neutral excitations, is often limited by the difficulty of separating the putative unconventional contributions from the phonon background. In practice, one needs either comparison with nonmagnetic reference compounds or performing a power-law fitting of the phononic $\kappa_{ph}(T)$ curve over a restricted temperature range. Both approaches are tricky and proliferate the discrepancies in interpreting the data.
This issue is exemplified in studies of $\alpha$-RuCl3, the most prominent Kitaev QSL candidate. Several key observations including complex magnetic field effect on $\kappa$ at low temperatures Leahy2017 , thermal Hall ($\kappa_{xy}$) effects Kasahara2018 , and quantum oscillations in $\kappa(H)$ Czajka2021 , were initially recognized as signatures of some QSL states. Subsequent studies challenged them by trying to claim that magnetically perturbed phonons can also produce these novel phenomena Hentrich2018 ; Lefrancois2022 ; Bruin2022 . This controversy further intensifies when Na2Co2TeO6 (NCTO) was found to exhibit some key experimental results similar to $\alpha$-RuCl3 Yao2020 ; Lin2021 ; Hong2021 ; Sanders2022 ; Yang2022 ; Gillig2023 ; Chen2024 ; Zhou2024 . While NCTO initially emerged as another potential Kitaev QSL material Liu2018 ; Sano2018 ; Liu2020 , detailed theoretical calculations indicate that the Kitaev interaction, while present, is not the dominant term in NCTO’s Hamiltonian Winter2022 . Taken together, that implies there might be nothing beyond trivial contributions to $\kappa$ in NCTO, and that conclusion might also be true for $\alpha$-RuCl3. In this context, it is important to get more detailed knowledge of the phonons in this material class.
@@
In this work, we focus on measuring the impact of out-of-plane ($\bf B$$\bf c$) magnetic fields on the in-plane ultrasound properties of NCTO. Our results demonstrate that ultrasound serves as a sensitive probe of magnetic transitions in NCTO. Very rich phenomena are revealed below its Néel temperature ($T_{N}=27$ K). By analyzing our results, we confirm the existence of an out-of-plane ferrimagnetic order in NCTO and provide evidence supporting a triple-Q ground state Kruger2023 ; Chen2021 . Furthermore, we explore the relationship between our ultrasound data and the reported thermal transport results under out-of-plane magnetic fields Yang2022 ; Hong2024 ; Li2023 , offering new insights into the interplay between lattice dynamics and magnetic excitations in this material and beyond.
@@ II Experimental Method
[…]
@@ III Results and Discussion
The temperature dependence of the relative changes in the sound velocity $\Delta v/v_{0}$ for the longitudinal acoustic (LA) mode is shown in Fig. 1(a). The overall $\Delta v/v_{0}(T)$ curve reveals the lattice hardening upon cooling, and a softening appears below the magnetic transition temperature $T_{N}$. The data at low temperature is enlarged in Fig. 1(b), plotted together with results taken with various ultrasound frequencies, to check for a possible frequency dependence. All these curves overlap almost perfectly. Three anomalies can be reliability resolved: the primary antiferromagnetic transition at $T_{N}=27$ K and two additional features at $T^{\prime}=16$ K and $T^{\prime\prime}=4$ K. Although this hierarchy of magnetic transitions below $T_{N}$ in NCTO has been previously documented Lefrancois2016 ; Bera2017 , it remains scarcely addressed. A recent comprehensive study combined spectroscopic and scattering techniques attributed this behavior to the interplay between two-dimensional and three-dimensional magnetic orders in NCTO Chen2021 .
@@
A set of field-dependent $\Delta v/v_{0}$ isotherms are presented in Fig. 2(a). Above 5 K, the $\Delta v/v_{0}(H)$ curves show moderate lattice softening in field without additional features up to $7.5$ T. The amplitude of changes in $\Delta v/v_{0}(H)$ increases considerably upon cooling. A clear dip in $\Delta v/v_{0}(H)$ can be resolved at 3 K. This dip gradually shifts to a lower field at lower temperatures. At the lowest temperature of 32 mK another broader dip at round 16 T was detected. These features are summarized in Fig. 2(b) in the phase diagram of NCTO for an out-of-plane magnetic field, together with the transitions observed in the magnetic susceptibility Zhang2024b . While all the ultrasound anomalies can find their correspondence in magnetization, some magnetic transitions reported in the literature leave no detectable fingerprint in our ultrasound data, including $H_{1}$ and $H_{0}$ above 3 K  Zhang2024b ; Bera2023 . This fact means that the acoustic modes in NCTO are selectively coupled to the spins. For the studied acoustic modes, the magnetoelastic coupling can be very small at some phase transitions in NCTO.
Figure 2: (a) Field dependence of $\Delta v/v_{0}$ at different temperatures. The curves are shifted vertically for clarity. Note the different rescale factors for each curve. The black arrows indicate the dips in some curves. (b) The gray circles and squares are phase boundaries of NCTO inferred from its magnetic susceptibility data Zhang2024b . The red stars are the positions of anomalies from Fig. 2(a). We follow the notation for the critical fields $H_{1}$ and $H_{2}$ from Ref. Zhang2024b . The dashed gray lines are guide-to-the eyes of the boundaries of the contour plot of $dM/dH$ data in Ref. Zhang2024b . Although there is a clear line below $H_{1}$ observable, it was not labeled as a critical field in Ref. Zhang2024b . Nevertheless, an electron spin resonance (ESR) work identified it also as a phase transition Bera2023 . Thus we use $H_{0}$ to name this line. The dashed black lines indicate the parameter space reached in Fig. 2(a).
Figure 3: Field dependence of (a) $\Delta v/v_{0}$ and (b) $\Delta\alpha$ at 1.8 K, measured after cooled in opposite field directions. The first field-sweep run was done after the sample was cooled in zero, or in small (<14<14 mT) remanent field of our magnet. The second and third measurements were done after cooling in a finite field applied along opposite directions.
@@
Recently, there have been some debates on whether the magnetic ground state of NCTO is an unusual triple-Q state or the conventional multi-domain zigzag order Chen2021 ; Yao2023 ; Jin2025 ; Zhang2023 ; Dhanasekhar2025 . To shed more light on this issue, we investigated field dependence of $\Delta v/v_{0}$ and $\Delta\alpha$ under different cooling protocols (Fig. 3). We observe a marked field-asymmetry tied to cooling conditions, switchable by modifying the field-cooling procedure above $T_{N}$. This provides clear evidence of an out-of-plane ferromagnetic moment, consistent with the reported out-of-plane ferrimagnetic behavior in NCTO Yao2020 . There are two mechanisms proposed for the ferrimagnism in NCTO, both invoking the canting of the in-plane magnetic moments. Some previous work suggested a triple-Q ground state in NCTO and several other Co-based quasi-2D honeycomb magnets Chen2021 ; Yao2023 ; Jin2025 ; Kruger2023 ; Francini2024 . Another scenario insists that many cobaltates manifest a zigzag ground state stabilized within an XXZ model Zhang2023 ; Zhang2024b ; Dhanasekhar2025 ; Winter2022 , and the out-of-plane ferrimagnetism is due to unbalanced domains. Note that the zero-field-cooled (ZFC) $\Delta v/v_{0}(H)$ and $\Delta\alpha(H)$ data perfectly match the field-cooled (FC) data of the fully-switched states. It is unlikely that the alignment of domains can be made by a small remanent field. Hence, a more natural explanation is provided by a domain-free triple-Q ground state of NCTO.
@@
To gain more insight, one set of additional $\Delta v/v_{0}(H)$ results at elevated temperatures all the way up to above $T_{N}$ are shown in Fig. 4(a). Results collected with different ultrasound frequencies and polarizations can be found in Supplemental Material SM . Obviously, the $\Delta v/v_{0}(H)$ curves are weakly field-suppressed and quite symmetric at temperatures higher than $T_{N}$. A butterfly-shaped hysteresis starts to grow at low temperatures. The hysteretic region and asymmetry increase as the temperature is lowered. Finally, the hysteresis quickly shrinks, leaving very asymmetric $\Delta v/v_{0}(H)$ curves below about 12 K.
Figure 4: (a) $\Delta v/v_{0}(H)$ isotherms measured in NCTO in opposite field directions. The curves are shifted vertically for clarity. (b) The temperature dependence of the integrated hysteretic region of $\Delta v/v_{0}(H)$ isotherms normalized to the maximum change of $\Delta v/v_{0}$ at each temperature. The integration of $\Delta v/v_{0}(H)$ hysteresis for data collected under different experimental settings (see text) are rescaled to show a common temperature behavior. The yellow stripe is a guild to the eye.
Figure 5: Field dependence of $\Delta v/v_{0}$ and $\Delta\alpha$ at selected temperatures: (a) 300 mK, (c) 625 mK, (e) 3 K, (g) 5 K, and (i) 15 K. The corresponding estimates of the change of phonon conductivity (see text) are plotted in red in (b), (d), (f), (h) and (j), respectively. Field dependence of the thermal conductivity $\Delta\kappa/\kappa_{0T}$ at similar temperatures reported in the literature is also shown Yang2022 ; Hong2024 ; Li2023 . For T=T=5 K, the ultrasound data of two frequencies were presented. (k) Field dependence of $\Delta\kappa/\kappa_{0T}$ and $\Delta v/v_{0}$ of $\alpha$-RuCl3 at comparable temperatures, with the field applied along the same in-plane crystallographic direction. Data are from Refs. Hauspurg2024 ; Czajka2021 .
[…]
Naively, one might anticipate stronger field hysteresis corresponding to more substantial reorganization of spin configurations. If NCTO has a XXZ ground state with domains developing below $T_{N}$, the spin-flip energy barrier should exhibit monotonic enhancement with decreasing temperature. Consequently, the observed hysteresis maximum occurring well below $T_{N}$ is an unconventional behavior. On the other hand, a recent Monte Carlo simulation work deduced a compensation point $T_{cp}$ at roughly $T_{N}/2$ for a sign reversion of the residual magnetization out of the honeycomb plane, if a triple-Q model is adopted for NCTO Francini2024 . Given that the energy difference between two magnetic configurations vanishes near 12 K, an applied magnetic field within this temperature regime becomes particularly effective in modifying the magnetic order. Taken together, our observations are better interpreted by a triple-Q ground state of NCTO.
@@
$\kappa\propto Cvl=Cv^{2}\tau\propto\frac{Cv^{2}}{\Gamma}~~,$ (1)
@@
$\Delta\kappa\propto\frac{C(v_{0}+\Delta v)^{2}}{\Gamma_{0}+\Delta\Gamma}-\frac{Cv_{0}^{2}}{\Gamma_{0}}~~.$ (2)
@@
$\frac{\Delta\kappa}{\kappa}\approx 2\times\frac{\Delta v}{v_{0}}-\frac{\Delta\Gamma}{\Gamma_{0}+\Delta\Gamma}~~,$ (3)
@@
$\Delta\alpha\equiv-\frac{20}{L}\log_{10}\frac{I}{I_{0}}~~,$ (4)
@@
$\frac{\Gamma_{0}}{\Gamma}=\frac{\tau}{\tau_{0}}\propto\frac{I}{I_{0}}=10^{-\Delta\alpha\times L/20}~~.$ (5)
@@
$\frac{\Delta\kappa}{\kappa}\approx 2\times\frac{\Delta v}{v_{0}}+10^{-\Delta\alpha\times L/20}-1~~.$ (6)
The estimated relative change in phonon thermal conductivity based on ultrasound data is shown in Fig. 5, by taking the sample thickness $L=0.4$ mm in Eq. (6). The effect of the out-of-plane magnetic field on the thermal transport of NCTO at nearly the same temperature reported in the literature is also sketched for comparison Yang2022 ; Hong2024 ; Li2023 . At temperatures below 1 K [Fig. 5(b) and (d)] and at around 15 K [Fig. 5(j)], the shape of the calculated and measured $\Delta\kappa(H)$ curves roughly matches, but the amplitude calculated from the phonon data is underestimated by an order of magnitude or more. The reason might be that only one acoustic mode at a single frequency is considered in this work, while all phonon modes contribute to phonon thermal transport. Actually, the $\Delta\alpha$ term is strongly frequency dependent SM . It is justified to claim that these results are consistently explained by the dominant phonon contribution of total thermal conductivity. However, as clearly demonstrated by Fig. 5(f) and (h), at intermediate low temperatures of 3 K and 5 K, while an impressively large field suppression of $\kappa$ is expected from the ultrasound data, the measured $\Delta\kappa(H)$ are relatively modest (overestimated by a factor of 2), and its shape does not follow the phonon expectation. Hybridization of phonons and magnons was proposed by Ref. Li2023 to explain the thermal Hall experimental results observed in NCTO. We argue that such hybridization might also be responsible for the discrepancy between real $\Delta\kappa(H)$ and the estimation from our ultrasound data in the intermediate low temperature range. If phonons and magnons exhibit synergistic transport behavior, the simplified scattering picture underlying Eq. 6 becomes inadequate. We expect that the comparison between ultrasound and thermal transport data can help delineate the temperature range where the hybridization effect prevails.
The literature $\Delta v/v_{0}(H)$ and $\Delta\kappa(H)$ curves of $\alpha$-RuCl3 at the similar temperature are plotted in Fig. 5(k) Hauspurg2024 ; Czajka2021 . Both show clear oscillatory structures above the two well-defined critical fields. This oscillatory structure in $\kappa(H)$ was initially attributed to some charge neutral magnetic excitations Czajka2021 , and was later challenged by a more conventional scenario of strongly scattered phonon transport channels Bruin2022 ; Lefrancois2023 . In the context of our results and discussion, the scattered phonon explanation and possibility of phonon-magnon hybridization should also be carefully considered when interpreting these novel $\kappa(H)$ data.